OpenStudy (anonymous):

19. 2x + 3y = 2 3x + 4y = 7

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (tkhunny):

Is there a specific method that is recommended?

OpenStudy (anonymous):

addition or subtraction

OpenStudy (tkhunny):

Still quite a few ways to go about it. Here's one such way... Stare at the coefficients on the x's. What is the least common multiple of 2 and 3?

OpenStudy (anonymous):

1

OpenStudy (tkhunny):

No, that's a divisor. We need a multiple.

OpenStudy (anonymous):

6

OpenStudy (tkhunny):

That's better. Multiply the first equation by 3 Multiply the second equation of 2 See if anything interesting results.

OpenStudy (anonymous):

3x+3y=15

OpenStudy (tkhunny):

Whence came that? That isn't anything. Is that the altered 1st equation, the altered 2nd equation, or something else?

OpenStudy (anonymous):

x+y=5

OpenStudy (anonymous):

I multiplied the first equation by 3 and the second by 2 and then subtractetd

OpenStudy (tkhunny):

STOP!! I don't know where you are going. Start here. 2x + 3y = 2 3x + 4y = 7 Multiply the 1st equation by 3 6x + 9y = 6 Multiply the 2nd equation by 2 6x + 8y = 14 You should get to that point.

OpenStudy (anonymous):

18x+27y=18 12x+16y=28

OpenStudy (tkhunny):

Why did you multiply them again? Just do it once.

OpenStudy (anonymous):

oh srry

OpenStudy (tkhunny):

After Multiplying the 1st equation by 3 and the 2nd equation by 2, you are left with these two equations. 6x + 9y = 6 6x + 8y = 14 It is time to move to the subtraction.

OpenStudy (anonymous):

y=-8

OpenStudy (tkhunny):

Almost magic, wasn't it? Okay, now how do you suggest we find a value for x?

OpenStudy (anonymous):

plug in y

OpenStudy (tkhunny):

Note: "plug in" Never say that. Please feel free to SUBSTITUTE the given value for y into any ORIGINAL equation.

OpenStudy (anonymous):

2x + 3(-8) = 2 3x + 4(-8) = 7 2x-24=2 3x-32=7 2x=26 x=13 3x=39 x=13

OpenStudy (tkhunny):

Fair enough. Notice how only one was necessary?

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!

OpenStudy (tkhunny):

No worries. It's what we do around here! :-)